OpenStudy (anonymous):

Rectangle A has an area of 4 - x2. Rectangle B has an area of x2 + 2x - 8. In simplest form, what is the ratio of the area of Rectangle A to the area of Rectangle B? Show your work.

jhonyy9 (jhonyy9):

x^2 +2x -8=(x+4)(x-2)

OpenStudy (mertsj):

\[\frac{4-x^2}{x^2+2x-8}=\frac{-1(x^2-4)}{x^2+2x-8}\]

jhonyy9 (jhonyy9):

4-x^2 =-(x^2 -4) =-(x-2)(x+2)

OpenStudy (mertsj):

\[\frac{-1(x-2)(x+2)}{(x+4)(x-2)}\]

OpenStudy (mertsj):

Cancel the common factor.